find the present value of $3000 comp monthly for 6% for 7 years.
I answered 3000/(1=.06/12)^84= 1973.20 or 2532.60.

Answer by checkley77(12569)

3000{/}(1{=}.06/12)^84= 1973.20 or 2532.60.
( ) incorrect modifiers.
3000(1+.06/12)^[7*12]=3000(1+.005)^84=3000(1.005)^84=3000*1.520=$4,561.11 ans.